OpenStudy (anonymous):

It is possible to make a usable thermometer out of water. Suppose you have a 500 mL glass bottle which is filled with water at 16 C. You put a 8 mm diameter straw into the water and seal with clay so that the level of the water in the straw is exactly at the top of the clay (assume that the thermal expansion of the bottle can be neglected). If you raise the temperature of the water to 45, how far up the straw does the water level move? Express your answer using three significant figures and cm.

OpenStudy (anonymous):

What I tried: I found that thermal expansion from 0-100C is -0.042? Is that right? Am I suppose to use that...? -0.042 x 500 x (45-16)= -609 mL \[\Pi*r*h^2\] \[\Pi*4^2*h=-609\] h=12.1cm Which was wrong.

OpenStudy (anonymous):

why are you getting negative answers? :P and thermal expansion is from 16 to 45 degrees..

OpenStudy (anonymous):

http://www.engineeringtoolbox.com/water-thermal-properties-d_162.html I used that, but I guess that was a dash, oops.

OpenStudy (anonymous):

Still, I took out the negative by the end because I knew that didn't make sense. I just don't know what constant I'm suppose to use for that number? Or is it not a constant at all? I'm confused.

OpenStudy (anonymous):

see for liquids. u have to use the volume coefficient of expansion.. so use this table http://www.engineeringtoolbox.com/cubical-expansion-coefficients-d_1262.html then use deltaV = (vol coeff).V.(delta T)

OpenStudy (anonymous):

Okay, so 0.000214 for water at Celsius. But I saw the foot note and noticed that "volumetric expansion coefficients for water at different temperatures" which threw me off. Should I just be using 0.000214?

OpenStudy (anonymous):

yea yea.. u can neglect the variation of the volume coefficient with temperature.

OpenStudy (compassionate):

Hi, I took the liberty of solving all your problems for you. http://lmgtfy.com/?q=what+is+the+answer+to+life%2C+the+universe%2C+and+everything ~Cheers

OpenStudy (anonymous):

@Compassionate wow.. that solved all my problems :O :O.. thank you :O

OpenStudy (compassionate):

That's what Ambassadors are for. Glad I could help.

OpenStudy (anonymous):

what would i do without you? :P

OpenStudy (compassionate):

Probably die.

OpenStudy (anonymous):

delta v = (V coeff)(V)(delta T) delta v = (0.000214)(500)(45-16) V= 31.03 cm?

OpenStudy (compassionate):

V = 42.

OpenStudy (anonymous):

I'm serious. I'm getting docked points for every wrong answer I input.

OpenStudy (anonymous):

that is delta v.. so increase in volume u are getting is 3.103 cc (check ur calculation) so this is the amount of water that will rise up in the straw.. now u can calculate the height right?

OpenStudy (anonymous):

now be care full with the units while u do the calculation!

OpenStudy (anonymous):

V= L*W*H 3.103 cm = 0.8 cm * 4 * H 3.103 = 32h h = 0.10 Uhhh I don't know if I did that right. I don't want to input it because I'm on my last guess before I get no credit.

OpenStudy (anonymous):

Oh wait, I used the radius twice... not sure why V = L * W * H 3.3103 = L * 0.8 * H I'm not sure what the length is, I suppose.

OpenStudy (anonymous):

its 3.103 centimeters CUBE.. 3.103 cm^3 = pi* r^2* h.. good u didn't enter that :D

OpenStudy (anonymous):

its a strawwwwwwwwwwwwwww.. not a cuboid :D

OpenStudy (anonymous):

3.103 = Pi*0.4^2*h 3.103 = 0.50h h= 6.22 cm? Ugh, I knew that. That's what I was trying to do before. I swear to god, I don't know why this is causing me so many problems. I just get flustered easy ;;

OpenStudy (anonymous):

thats ok .. lemme check that wait huh!

OpenStudy (anonymous):

i am getting 6.18 cm

OpenStudy (anonymous):

I think I just rounded a step early instead of pulling the whole number. It said it was exactly 6.06, but accepted that answer. Thanks for the help :3

OpenStudy (anonymous):

no problem.. you had it all along.. i just gave u a small nudge to reach to the final answer :D